OpenStudy (anonymous):

Solve using the elimination method. If the system has no solution or an infinite number of solutions, state this. 10x – y = 12.5 3x + y = 7

OpenStudy (anonymous):

do you add the two equations to cancel out the y

Parth (parthkohli):

Yes, you do!

OpenStudy (anonymous):

10x - y = 12.5 3x + y = 7 13x = 19.5 x = 1.5 10(1.5) - y = 12.5 15 - y = 12.5 -y = .83333 y = - .833333

OpenStudy (phi):

good up to 15 - y = 12.5 or 15= 12.5+y 15-12.4=y

OpenStudy (phi):

*15-12.5= 2.5

OpenStudy (anonymous):

10(1.5) - -.833 = 12.5 15 + .8333 = 12.5 15.8333 = 12.5 3(1.5) + -.833 = 7 4.5 - .833 = 7 3.667 = 7 is this right

OpenStudy (phi):

of course not. 15.8333 does not equal 12.5 see post above for your mistake

OpenStudy (anonymous):

i am writing this down so might take a min ok i am looking now

OpenStudy (anonymous):

why did you rearrange the problem? it it to get y by itsself

OpenStudy (phi):

habit. I don't like -y= stuff . It does not matter until you do < problems instead of = if you like: 15 - y = 12.5 -y= 12.5-15 -y= -2.5 y= 2.5 (multiply both sides by -1)

OpenStudy (anonymous):

all you do is substitute then right

OpenStudy (phi):

yes (to check the answer) and it should work this time.

OpenStudy (anonymous):

ty for your help